OpenStudy (anonymous):

Given the exponential equation 3x = 243, what is the logarithmic form of the equation in base 10? x = log base 10 of 243, all over log base 10 of 3 x = log base 10 of 3, all over log base 10 of 243 x = log base 2 of 3, all over log base 2 of 243 x = log base 2 of 243 all over log base 2 of 3

OpenStudy (solomonzelman):

\(\large\color{black}{ \rm 3^{x}=243 }\) USE, \(\large\color{black}{ \rm a^{b}=c~~~~~~->~~~~~~~log_ac=b }\)

OpenStudy (anonymous):

x = log base 10 of 243, all over log base 10 of 3 x = log base 10 of 3, all over log base 10 of 243 x = log base 2 of 3, all over log base 2 of 243 x = log base 2 of 243 all over log base 2 of 3

OpenStudy (sweetburger):

no its none of the answer choices you gave us

OpenStudy (sweetburger):

\[a^y=x\] is equivalent to \[\log_{a} x=y\] so you gave us i believe 3^x=243 so 3=a, x=y, and 243=x

OpenStudy (sweetburger):

you may want to rewrite your question as 3^y=243 so that y=y and there is no confusion